OpenStudy (anonymous):

Write an equation of the line containing the specified point and parallel to the indicated line. Please show work. (-7,0),5x+2y=6

OpenStudy (anonymous):

First you will need to find the slope of the given line. Since we are trying to generate a parallel line we will want to have the same slope. Solve the given equation for y to have it in the from y = mx + b, and m will be your slope. Then you can use that slope and the point (-7,0) to find the line using the point-slope form \[y-y _{1}=m(x-x _{1})\]. Where -7 is the \[x _{1}\] and 0 is the \[y _{1}\].

OpenStudy (anonymous):

Get the slope by rewriting the equation of that line in the form \(y=mx+c\), which gives \[y=-\frac{5}{2}x + 3\ \: => m = -\frac{5}{2}\] Use the equation of the line formula to get the new equation: \[y - y_{1} = m(x - x_{1}) \: ; \: x_{1} = -7 \: ; \: y_{1} = 0 \] Which yields \[y = -\frac{5}{2}x - \frac{35}{2}\]